1. Prove that n + log n = O(n) by showing that there exists a constant c > 0 such that n + log n ≤ cn.
1
Expert's answer
2019-02-25T08:23:03-0500
"n+\\log \u2061n\\le n+\\log\u2061(1+n+n^2\/2!+n^3\/3!)\\le""n+\log( e^n) =n+n=2n,\quad n\in \mathbb{N}"